OpenStudy (anonymous):

A farmer plans to use 21 m of fencing to enclose a rectangular pen having an area of 55 m². Only three sides of the pen need fencing because of an existing wall. Find the dimensions of the pen.

OpenStudy (anonymous):

For this problem you need to deal with both perimeter and area so let's start with those formulas If the length is denoted by l and the width by w: P = 2l + 2w A = l Now we only are concerned with three sides of the pen when dealing with perimeter so using the 21m of material he has, we see that 21 = l + 2w (since the other length is using the existing wall. We want the area to be 55m^2 so plugging this in we have 55 = lw Solving the perimeter equation for l and substituting, we get\[55=lw=(21-2w)w=21w-2w^{2}\]\[-2w^{2}+21w=55\] We can now solve this quadratic by moving the 55 to the left and factoring \[-2w^{2}+21w-55=0\]\[2w^{2}-21w+55=0\](2w - 11)(w - 5)=0 Solving each of these factors for w will give us two posible values for the width which we can plug into the perimeter equation to get the associated lengths. I leave the rest up to you...

OpenStudy (anonymous):

5 by 11 (by eye)

OpenStudy (anonymous):

That's only one of the answers @estudier. Since it's a quadratic, there are 2 solutions.

OpenStudy (anonymous):

5.5 m by 10 m or 5 m by 11 m

OpenStudy (anonymous):

The other is probably something daft, like a negative (guessing)

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

What's the max area you can enclose?

OpenStudy (anonymous):

@nany - You are correct! Outstanding work! @estudier - No, the other solution is not daft. Many of these area/perimeter problems have to viable solutions. One is usually nice numbers like 5 x 11 and the other is less nice because you're reallocating "blocks" from the side and redistribuiing them to the length. Perhaps, @estudier, you should work out the problem using the corret method prior to assuming answers and giving false information.

OpenStudy (anonymous):

Thank You So Much For Your Help !!

OpenStudy (anonymous):

A = x(21-2x) so dA = 21-4x = 0 for max A --> x=21/4 and max A = 55 and an eighth. @mtbender74, I gave no false information (My first answer I said was by eye and the second I said was a guess, I apologize if you misunderstood me).